# Robust Understanding of Human-Robot Social Interactions through Multimodal Distillation

### v1

@@ Abstract.
The need for social robots and agents to interact and assist humans is growing steadily. To be able to successfully interact with humans, they need to understand and analyse socially interactive scenes from their (robot’s) perspective. Works that model social situations between humans and agents are few; and even those existing ones are often too computationally intensive to be suitable for deployment in real time or on real world scenarios with limited available information. We propose a robust knowledge distillation framework that models social interactions through various multimodal cues, yet is robust against incomplete and noisy information during inference. Our teacher model is trained with multimodal input (body, face and hand gestures, gaze, raw images) that transfers knowledge to a student model that relies solely on body pose. Extensive experiments on two publicly available human-robot interaction datasets demonstrate that the our student model achieves an average accuracy gain of 14.75% over relevant baselines on multiple downstream social understanding task even with up to 51% of its input being corrupted. The student model is highly efficient: it is $<1$% in size of the teacher model in terms of parameters and uses $\sim 0.5$‰ FLOPs of that in the teacher model. Our code will be made public during publication.
@@
ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing HCI theory, concepts and models
Figure 1. Our aim is to build a lightweight and robust model to understand human-robot social interactions. A novel teacher-student learning framework is built to achieve this, where the teacher model uses multimodal cues (body pose, face and hand gestures, and gaze) and raw RGB frames to learn a rich representation of the social scenario. Through knowledge distillation, the teacher model is able to transfers the knowledge to a lightweight student model that relies solely on body pose features. The student model is robust against incomplete and noisy inputs.
@@ 1. Introduction
Recent application advances in social robotics and virtual agents include personal care, navigation, and user interaction (Zhou et al., 2022). A central challenge in social AI and human-robot interaction (HRI) is enabling the systems to understand and forecast human social behaviours from an egocentric perspective, such as intent to interact, attitudes, and social actions (Bian et al., 2025). This capability is particularly critical during the early stages of interaction, where understanding the user’s engagement intent allows the system to proactively prepare for smooth and natural interactions. Early prediction enhances user experience and supports adaptive, context-aware responses (Avelino et al., 2021). However, deploying such predictive models in real-world settings remains challenging due to constraints in computational resources, input quality, and the need for real-time processing.
@@
Understanding human social behaviours relies on multiple non-verbal signals (Urakami and Seaborn, 2023), including body, face and hand gestures, gaze and raw images, each providing unique and complementary information into social interactions (Yumak and Magnenat-Thalmann, 2015). Despite progress in understanding multimodal nonverbal cues, several challenges remain in using them for broader social understanding in real-world deployment. (1) Multimodal Integration: Social understanding needs fusing information from diverse modalities that can sometimes be inconsistent. Models need to integrate information and differences across modalities to extract coherent semantic representations. (2) Real-World Challenges: Egocentric perspectives introduce challenges such as limited field of view, camera motion, occlusions, and feature estimation errors, which degrades the reliability of social understanding. (3) Computational Efficiency: Using multiple modalities requires computationally intensive models and more feature estimation modules, posing difficulties for real-time inference on resource-constrained systems, such as social robots. Addressing these challenges needs a lightweight yet robust framework that can integrate multimodal cues under noisy conditions while supporting real-time social understanding.
@@
To tackle these challenges, we propose a knowledge distillation framework that enables a lightweight student model to infer rich social cues from incomplete body pose input alone and jointly forecast three key aspects of social understanding: interaction intentions, attitudes, and social actions, as shown in Fig.1. Our approach consists of two stages: first, we train a teacher model with full-modal inputs, including body, face and hand gestures, gaze and raw images, to learn and output high-quality social representations. Then, we distil this knowledge into a lightweight student model that relies solely on body pose while incorporating spatio-temporal noise to simulate real-world data degradation. Instead of only learning from downstream task labels, the student also learns to infer multimodal and clean social cues through knowledge distillation, ensuring robust performance under noisy conditions. The lightweight design of the student model improves computational efficiency.
@@
In summary, our work enhances the robustness and efficiency of jointly forecasting users’ intent to interact, attitudes, and social actions from an egocentric perspective through the contributions:
@@
    Robustness to real-world noises: To enhance robustness against noisy and incomplete input, we introduce spatiotemporal corruption for the student model. With knowledge distillation, when 51% of spatiotemporal information is lost the student maintains an average accuracy of 81.35% on JPL-Social (Bian et al., 2025) (14.68% improvement over independent training) and 59.05% on HARPER (Avogaro et al., 2024) (14.82% improvement).
@@
    Efficiency and deployment readiness: The proposed student model SocialEgoMobile demonstrates high computational efficiency, requiring only 1% of model parameters and 0̃.5‰ of the floating-point operations (FLOPs) compared to the teacher model.
@@ 2.1. Social Intent Prediction
Egocentric social intent prediction is crucial for socially intelligent agents to establish natural, harmonious, and efficient interactions with users. A precise understanding of the user’s social behaviour enables agents to timely adjust response strategies to meet user needs. Current research primarily models social intent as a binary classification task, focusing on predicting whether a user intends to interact with an intelligent agent. Approaches include analyzing motor cues such as body orientation, movement direction, and speed (Abbate et al., 2024), targeting users interested in interaction using visual and audible features (Pourmehr et al., 2017), and using gaze information in virtual environments to infer interaction intent (David-John et al., 2021). However, these methods are limited to distinguishing between ‘interact’ and ‘not interact’, failing to capture the richness and complexity of users’ social intentions.
@@ 2.2. Non-Verbal Cues in Multimodal HRI
In Human-Robot Interaction (HRI), accurately interpreting the user’s identity, social intent, behaviours, and communication requires integrating information from multiple modalities (Yumak and Magnenat-Thalmann, 2015). Meanwhile, non-verbal cues play a crucial role in understanding user intentions and needs (Urakami and Seaborn, 2023). For instance, while detecting sarcasm through text is challenging during a conversation, recognizing subtle body movements such as shoulder shrugs or lip curls allows humans to accurately identify sarcasm (Castro et al., 2019). Consequently, capturing and interpreting mutlimodal non-verbal cues in HRI is essential for assessing the users. Recent research has explored various strategies for multi-modal fusion to enhance the accuracy of interaction models. For example, combining RGB images, facial expressions, and distance information can effectively assess user engagement (Ben-Youssef et al., 2019) and interest (Foster et al., 2019) when interacting with service robots in public spaces. Facial expression and gaze information can be used to infer whether a user requires assistance in collaborative tasks (Wilson et al., 2022). To further improve robustness in real-world scenarios, researchers tried to model temporal information of non-verbal signals. Recurrent models have been used to model gestures and RGB video streams for identifying user action commands (Liu et al., 2018), and to detect user emotions during collaboration by analyzing body movements and facial expressions (Páez and González, 2022). Pose skeletons and RGB information have been fused to predict interaction actions in single-user (Islam and Iqbal, 2020) and multi-user scenarios (Yasar et al., 2024). Non-verbal modalities including body, face, hand gestures, gaze and raw images provide rich contextual information that enhances a robot’s ability to accurately understand to users’ intentions, attitudes, and actions, improving the effectiveness of HRI systems.
Research gap. Despite the progress in social understanding using deep learning models, existing approaches often rely on multimodal inputs. However, processing more modal information increases computational cost (Liang et al., 2024) and the deep models are sensitive to missing or noisy information (Drenkow et al., 2021), making them impractical for deployment on resource-constrained devices and real-world scenarios. While knowledge distillation has been explored to compress large models into lighter versions, most prior works focus on preserving accuracy under ideal conditions, overlooking the need to integrate multimodal information using a single modality and maintain robustness in environments with incomplete or degraded inputs (Arani et al., 2021). To address these limitations, we propose a knowledge distillation framework with a lightweight, body pose-only student model that learns from a full-modality teacher model through a robust distillation, ensuring both computational efficiency and resilience to temporal and spatial noise.
@@ 3. Methodology
Figure 2. Our knowledge distillation framework uses SocialC3D as the teacher model, which fuses raw images, body, face, hand gestures, and gaze. Each modality is processed by a ResNet (He et al., 2016) and integrated via lateral connections and late fusion, producing a high-quality social representation for downstream tasks. The lightweight student model, SocialEgoMobile, uses only corrupted body pose. It consists of a two-layer GAT (Veličković et al., 2018) and a single-layer Bi-LSTM (Graves and Schmidhuber, 2005) to extract social representations. The framework distillates knowledge from the teacher model by maximising the mutual information of the social representations output by the teacher and student model.
@@ 3.1. Problem Definition
Given a sequence of egocentric videos captured by an intelligent social agent, our aim is to forecast the user’s social intent and behaviours in real time based only on the early stages of interaction (e.g., the first second). Specifically, we forecast the user’s intent to interact, attitude, and social actions. To achieve this, we propose a knowledge distillation framework where a lightweight student model is trained to rely on body pose features, while a teacher model uses full multimodal inputs (body, face, hand gesture, gaze and raw images), as shown in Fig.2. Body pose features, facial keypoints, and hand keypoints are extracted from the video using Alphapose (Fang et al., 2022) and gaze features are extracted using MCGaze (Guan et al., 2023). Formally, we define the input signals at step $t$ as follows:
@@
(1) $R_{teacher}=F_{teacher}(\boldsymbol{x}_{t:T}^{\text{Pose}},\boldsymbol{x}_{t:T}^{\text{Image}},\boldsymbol{x}_{t:T}^{\text{Face}},\boldsymbol{x}_{t:T}^{\text{Hand}},\boldsymbol{x}_{t:T}^{\text{Gaze}})$
where $R_{t}$ represents the high-level social representation extracted by the teacher model. To ensure efficiency and robustness, we introduce a lightweight student model $F_{student}$, which relies solely on body pose features $\boldsymbol{x}_{t}^{\text{Pose}}$:
(2) $R_{student}=F_{student}(\boldsymbol{x}_{t:T}^{\text{Pose}})$
where $R_{student}$ represents the social representation extracted by the student model.
For the downstream tasks of forecasting user’s intent to interact, attitude, and social actions, the social representations $R$ extracted by the teacher or student model are processed by the Chain hierarchical classifier (Bian et al., 2025), which simulates how human understand and forecast social signals, predicting the results of three downstrem tasks jointly:
@@ 3.2. Teacher Model
To effectively integrate multiple modalities, we employ a teacher model, SocialC3D which is based on PoseC3D (Duan et al., 2022), a spatio-temporal feature extraction framework using 3D convolutional neural networks (3D-CNN). PoseC3D converts sequences of 2D pose keypoints into 3D heatmap stacks, according to their coordinates and confidence scores. The heatmaps are converted as follows:
@@
where $\sigma$ is the Gaussian variance of the heat map; ($x$, $y$) are the coordinates of the heatmap; ($x_{i}$, $y_{i}$) and $c_{i}$ are the coordinates and confidence scores of the current keypoint $i$. This approach effectively captures spatio-temporal dependencies and exhibits greater robustness to noise than traditional skeleton-based methods. Additionally, the lateral connections between the RGB and Pose pathway fuse image and pose information, enhancing multimodal learning.
Based on this, we introduce SocialC3D to incorporate more modal information. SocialC3D transforms additional modalities, facial expressions, hand gestures, and gaze, into 3D heatmap stacks as input, which are processed using the same ResNet (He et al., 2016) as the body pose stream. For gaze estimation, we use the centre of the head bounding box as the beginning and calculate the end based on the gaze direction and the diagonal length of head bounding box. SocialC3D incorporate lateral connections to propagate intermediate representations across modalities at multiple stages, enabling deeper feature integration.
@@ 3.3. Student Model
The student model takes only body pose keypoints as input. Compared to raw visual information, body language provides a more direct and interpretable signal for understanding social interactions. Additionally, body posture inherently contain some facial and hand dynamics, allowing the model to integrate multimodal cues through knowledge distillation. Moreover, body pose features, as high-level abstractions, support lighter spatio-temporal analysis models and pose estimation, enhancing the student model’s suitability for resource-constrained settings.
To reduce computational complexity while maintaining strong predictive performance, we developed SocialEgoMobile as our student model. To reduce the risk of overly deep networks amplifying noise and hindering useful information extraction, we designed a streamlined architecture consisting of a two-layer graphical attention network (GAT) (Veličković et al., 2018) and a single-layer bi-directional long- and short-term memory (Bi-LSTM) (Graves and Schmidhuber, 2005). By dynamically adjusting the attentional weights between neighbouring joints, GAT can effectively capture the spatial relationships between key points of the body pose skeleton graph, thus making the model more focused on useful information and limiting the noise. Meanwhile, Bi-LSTM enhances temporal modelling by processing sequences forward and backward with forgetting gates that can selectively filter noise. This lightweight yet effective design maintains stable performance under noisy and incomplete input conditions, thus enhancing robustness and making it suitable for real-time deployment in resource-limited environments.
@@ 3.4. Knowledge Distillation
To effectively transfer knowledge of the user’s social behaviour from the teacher model to the student model, we adopt a feature-based knowledge distillation method. The student model learns not only from the ground truth labels of the downstream tasks but also by minimizing the discrepancy between its high-level social representations and those of the teacher model. We employ InfoNCE (Information Noise Contrastive Estimation) (Oord et al., 2018) to estimate and maximize the mutual information between teachers’ and students’ social representations. Specifically, InfoNCE enhances similarity between social representation of teacher and student models of the same input while reducing similarity between different inputs. During training, the loss function of the student model combines the mutual information loss from InfoNCE with the classification loss from the downstream tasks, weighted for back propagation. The loss of InforNCE is defined as:
@@
where $x$ is the teacher model’s social representation for a given input, $x^{+}$ is the representation from the student model for the same input, and $x_{j}$ is the representation from the student model for different inputs. The similarity function $f(x,x^{\prime})$ is defined using cosine similarity as follows:
@@
By knowledge distillation, the student model can still learn high-level social knowledge from the teacher model with limited modality and corrupted inputs. Furthermore, we introduce spatio-temporal corruption in the student model’s input, enhancing its ability to integrate multimodal social cues while maintaining robustness to noise and missing data.

### v2

@@ Abstract.
There is a growing need for social robots and intelligent agents that can effectively interact with and support users. For the interactions to be seamless, the agents need to analyse social scenes and behavioural cues from their (robot’s) perspective. Works that model human-agent interactions in social situations are few; and even those existing ones are computationally too intensive to be deployed in real time or perform poorly in real-world scenarios when only limited information is available. We propose a knowledge distillation framework that models social interactions through various multimodal cues, and yet is robust against incomplete and noisy information during inference. We train a teacher model with multimodal input (body, face and hand gestures, gaze, raw images) that transfers knowledge to a student model which relies solely on body pose. Extensive experiments on two publicly available human-robot interaction datasets demonstrate that our student model achieves an average accuracy gain of 14.75% over competitive baselines on multiple downstream social understanding tasks, even with up to 51% of its input being corrupted. The student model is also highly efficient - less than 11% in size of the teacher model in terms of parameters and its latency is 11.9% of the teacher model. Our code and related data are available at github.com/biantongfei/SocialEgoMobile.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755463ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing HCI theory, concepts and models
Figure 1. We aim to build a lightweight and robust model to understand human-robot social interactions. A novel knowledge distillation framework is built, where the teacher model uses multimodal cues (body pose, face and hand gestures, gaze and raw images) to learn a rich social representation. Through distillation, the lightweight student model that relies solely on corrupted body pose features learns multimodal social knowledge and is robust against incomplete and noisy inputs.
@@ 1. Introduction
Recent application advances in social robotics and agents include personal care, navigation, and user interaction (Zhou et al., 2022; Zachiotis et al., 2018; Blair et al., 2025). A central challenge in social AI and human-robot interaction (HRI) is enabling the systems to understand and forecast human social behaviours from an egocentric perspective, such as intent to interact, attitudes, and social actions (Bian et al., 2025). This capability is particularly critical during the early stages of interaction, where understanding the user’s interaction intent allows the system to proactively prepare for natural and timely interactions. Early prediction enhances user experience and supports adaptive, context-aware responses (Avelino et al., 2021). However, deploying such predictive models in real-world settings remains challenging due to constraints in computational resources, input quality, and the need for real-time processing.
@@
Understanding human social behaviours relies on multiple non-verbal signals (Urakami and Seaborn, 2023), each providing unique and complementary information into social interactions (Yumak and Magnenat-Thalmann, 2015). Despite progress in modelling multimodal cues, several challenges remain in using them for social understanding in real-world deployment. (1) Multimodal Integration: Social understanding needs fusing information from diverse modalities that sometimes are inconsistent. Models need to integrate information and differences across modalities to extract coherent semantic representations. (2) Real-World Challenges: Egocentric perspectives introduce challenges such as limited field of view, camera motion and occlusions, which degrade the reliability of social understanding. (3) Computational Efficiency: Using multiple modalities requires computationally intensive models and more feature estimation modules, posing difficulties for real-time inference on resource-constrained systems, such as social robots.
@@
To tackle these challenges, we propose a knowledge distillation framework that enables a lightweight student model to infer rich social cues from corrupted body pose input and jointly forecast three key aspects of social understanding: intent to interact, attitudes, and social actions, as shown in Fig.1. Our approach consists of two stages: first, we train a teacher model with multimodal inputs, including body, face and hand gestures, gaze and raw images, to learn high-quality social representations. Then, we distil this knowledge into a lightweight student model that relies solely on body pose while containing spatiotemporal noise to simulate real-world data degradation, ensuring robust performance under noisy conditions. Body pose features are crucial cues for social understanding (Bian et al., 2025) and rely only on lightweight feature extraction and analysis models.
@@
In summary, our work enhances the robustness and efficiency of egocentric social understanding through the contributions:
@@
    Robustness to noises: To enhance robustness against noisy and incomplete input, we introduce spatiotemporal corruption for the student model during distillation. When 51% of spatiotemporal information is lost, the student maintains an average accuracy of 81.35% on JPL-Social (Bian et al., 2025) (14.68% improvement over independent training) and 59.05% on HARPER (Avogaro et al., 2024) (14.82% improvement).
@@
    Efficiency and deployment readiness: The proposed student model SocialEgoMobile demonstrates high computational efficiency, requiring only 1% of model parameters and 11.9% latency compared to the teacher model SocialC3D.
@@ 2.1. Social Intent Prediction
Egocentric social intent prediction is crucial for social robots to establish natural, harmonious, and efficient interactions with users. A precise understanding of the user’s social behaviour enables robots to timely adjust response strategies to meet user needs. Current research primarily models social intent as a binary classification task, focusing on predicting whether a user intends to interact with an intelligent agent. Approaches include analyzing motor cues such as body orientation, movement direction, and speed (Abbate et al., 2024), targeting users interested in interaction using visual and audible features (Pourmehr et al., 2017), and using gaze information in virtual environments to infer interaction intent (David-John et al., 2021). However, these methods are limited to distinguishing between ‘interact’ and ‘not interact’, failing to capture the richness and complexity of users’ social intentions.
@@ 2.2. Non-Verbal Cues in Multimodal HRI
In Human-Robot Interaction (HRI), accurately interpreting the user’s identity, social intent and behaviours requires integrating multimodal cues (Yumak and Magnenat-Thalmann, 2015). Meanwhile, non-verbal cues play a crucial role in social understanding (Urakami and Seaborn, 2023). For instance, while detecting sarcasm through text is challenging, recognizing subtle body movements such as shoulder shrugs allows humans to accurately identify sarcasm (Castro et al., 2019). Consequently, capturing and interpreting mutlimodal non-verbal cues in HRI is essential for assessing the users. Recent research has explored various strategies for multi-modal fusion to enhance the accuracy of interaction models. For example, combining RGB images, facial expressions, and distance information can effectively assess user engagement (Ben-Youssef et al., 2019) and interest (Foster et al., 2019) when interacting with service robots in public spaces. Facial expression and gaze information can be used to infer whether a user requires assistance in collaborative tasks (Wilson et al., 2022). To further improve robustness in real-world scenarios, researchers tried to model temporal information of non-verbal signals. Recurrent models have been used to model gestures and RGB video streams for identifying user action commands (Liu et al., 2018), and to detect user emotions during collaboration by analyzing body movements and facial expressions (Páez and González, 2022). Pose skeletons and RGB information have been fused to predict interaction actions in single-user (Islam and Iqbal, 2020) and multi-user scenarios (Yasar et al., 2024). Non-verbal modalities including body, face, hand gestures, gaze and raw images provide rich contextual information that enhances a robot’s ability to accurately understand to users’ intentions, attitudes, and actions, improving the effectiveness of HRI systems.
Research gap. While deep learning models using multimodal cues have advanced social understanding, their high computational demands (Liang et al., 2024) and sensitivity to noisy or incomplete inputs (Drenkow et al., 2021) limit deployment on resource-constrained platforms. Existing knowledge distillation approaches primarily aim to retain accuracy under ideal conditions (Arani et al., 2021), neglecting robustness and integrating multimodal information using a single modality. To overcome these issues, we propose a robust distillation framework where a lightweight, body pose-only student model learns from a multimodal teacher, achieving efficient, resilient, and accurate social understanding.
@@ 3. Methodology
Figure 2. Our knowledge distillation framework uses SocialC3D as the teacher model, which fuses raw images, body, face, hand gestures, and gaze information, producing a multimodal social representation. The lightweight student model, SocialEgoMobile, uses only body pose features to output social representations. The framework maximises the mutual information of the social representations of the teacher and student model, to transfer social knowledge.
@@ 3.1. Problem Definition
Given a sequence of egocentric videos captured by a social robot, we aim to forecast the user’s social intent, attitude and action in real-time by observing the early stage (the first second). So we propose a knowledge distillation framework where a lightweight student model is trained to rely on body pose features, while a teacher model uses multimodal inputs (body, face, hand gesture, gaze and raw images), as shown in Fig.2. Body pose features, facial keypoints, and hand keypoints are extracted from the video using Alphapose (Fang et al., 2022) and gaze features are extracted using MCGaze (Guan et al., 2023). Formally, we define the input signals at step $t$ as follows:
@@
(1) $R_{teacher}=F_{teacher}(\boldsymbol{x}_{t}^{\text{Pose}},\boldsymbol{x}_{t}^{\text{Image}},\boldsymbol{x}_{t}^{\text{Face}},\boldsymbol{x}_{t}^{\text{Hand}},\boldsymbol{x}_{t}^{\text{Gaze}}),t\in[0,T]$
where $R_{teacher}$ represents the high-level social representation outputed by the teacher model and $T$ represents the length of the input sequence. The lightweight student model $F_{student}$ relies solely on body pose features $\boldsymbol{x}_{t}^{\text{Pose}}$:
(2) $R_{student}=F_{student}(\boldsymbol{x}_{t}^{\text{Pose}}),t\in[0,T]$
where $R_{student}$ represents the social representation outputed by the student model.
For the subtasks of forecasting user’s intent to interact, attitude, and social actions, the social representations $R$ extracted by the teacher or student model are processed by a Chain hierarchical classifier (Bian et al., 2025), which simulates how human infer others’ social intents, predicting the results of three downstream tasks jointly:
@@ 3.2. Teacher Model
To integrate multiple modalities, we employ a teacher model, SocialC3D which is based on PoseC3D (Duan et al., 2022), a spatio-temporal feature extraction framework using 3D convolutional neural networks (3D-CNN). PoseC3D converts sequences of 2D pose keypoints into 3D heatmap stacks, according to their coordinates and confidence scores. The heatmaps are converted as follows:
@@
where $\sigma$ is the Gaussian standard deviation of the heatmap; ($x$, $y$) are the position in the heatmap; ($x_{i}$, $y_{i}$) and $c_{i}$ are the coordinates and confidence scores of the keypoint $i$. This approach exhibits greater robustness to pose estimation error than skeleton-based methods (Duan et al., 2022). Also, the lateral connections between the RGB and Pose pathway fuse information on multiple levels.
Based on this, we introduce SocialC3D to incorporate additional modalities, including facial expressions, hand gestures, and gaze, into 3D heatmap stacks as input, which are processed using the same ResNet structure (He et al., 2016) as the body pose pathway. For gaze vector, we use the centre of the head bounding box as the beginning and calculate the end based on the gaze direction and the diagonal length of head bounding box. SocialC3D incorporate lateral connections to propagate intermediate representations across modalities at multiple stages, enabling deeper feature integration. After the ResNets, the outputs of different modalities are concatenated and passed through a linear layer to produce a 16-dimensional multimodal social representation.
@@ 3.3. Student Model
The student model takes only body pose features as input. Compared to raw visual information, body language provides a more direct and interpretable cue for social understanding and body pose is one of the most effective modalities. Additionally, body posture inherently contains some facial and hand dynamics, allowing the model to integrate multimodal cues through knowledge distillation. Moreover, body pose, as a high-level abstraction, supports lighter analysis models and pose estimation methods, enhancing the student model’s suitability for resource-constrained settings.
To achieve a balance between performance, robustness to noise and computational efficiency, we developed SocialEgoMobile as our student model, consisting of a two-layer graphical attention network (GAT) (Veličković et al., 2018) and a single-layer bi-directional Long Short Term Memory (Bi-LSTM) (Graves and Schmidhuber, 2005). By dynamically adjusting the attentional weights between neighbouring joints, GAT can capture the spatial relationships of the body pose skeleton for each frame, focusing on useful information and limiting the noise. Meanwhile, Bi-LSTM captures the temporal dynamic of the input sequence forward and backward and filters noise using forgetting gates. The hidden dimension of the GATs is 16, and that of the Bi-LSTM is 128. The outputs are projected into a 16-dimensional social representation via a linear layer. This lightweight yet effective design maintains stable performance under noisy input conditions, thus enhancing robustness and making it suitable for deployment in resource-limited environments.
@@ 3.4. Knowledge Distillation
To transfer social knowledge from the teacher model to the student model, we adopt a feature-based distillation method, InfoNCE (Information Noise Contrastive Estimation) (Oord et al., 2018). The student learns from the ground truth labels of the downstream tasks and the difference between its social representation and that of the teacher. InfoNCE estimates and maximises its mutual information by enhancing the similarity between the social representations of teacher and student models of the same video clip while reducing the similarity between different ones. The distillation loss is defined as:
@@
where $x$ is the teacher’s social representation for a given video clip, $x^{+}$ is the representation from the studen for the same clip, and $x_{j}$ is the representation from the student for other clips. The similarity function $f(x,x^{\prime})$ is defined using cosine similarity as follows: